OpenStudy (anonymous):

Use the following function. log4(x + 3) = 2 To solve the equation, first: Select one: a. Re-write the equation using the definition of a logarithm. b. Divide both sides by 2. c. Take the log of each side. d. None of the above.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

I got A?

OpenStudy (kc_kennylau):

What would you do to solve it?

OpenStudy (kc_kennylau):

Yep I would also do A first

OpenStudy (anonymous):

\[\large \log_b(\diamondsuit)=\clubsuit \iff b^{\clubsuit}=\diamondsuit\]

OpenStudy (anonymous):

should I go with A

OpenStudy (bradely):

yes

OpenStudy (bradely):

log4(x + 3) = 2 x+3 =4^2